OpenStudy (anonymous):

Write 88.2% as a decimal i know the answer i just need the explination

jimthompson5910 (jim_thompson5910):

tell me what you got for the answer

OpenStudy (anonymous):

0.882

jimthompson5910 (jim_thompson5910):

you get that by dividing 88.2 by 100 88.2/100 = 0.882 or you can just move the decimal point 2 spots to the left

jimthompson5910 (jim_thompson5910):

so in general x% = x/100

OpenStudy (amistre64):

the word "percent" refers to a division by 100 if we take a number: n 1% of n is equal to: n/100 a decimal refers to a number between 0 and 1, so if we let n=1, then 1/100 = 1% if you want 88,.2 of those ... then 88.2*1/100